# Red Queen: Safeguarding Large Language Models against Concealed Multi-Turn Jailbreaking

### v1

@@ Abstract
Content Warning: This paper contains examples of harmful language and plans.
The rapid progress of Large Language Models (LLMs) has opened up new opportunities across various domains and applications; yet it also presents challenges related to potential misuse. To mitigate such risks, red teaming has been employed as a proactive security measure to probe language models for harmful outputs via jailbreak attacks. However, current jailbreak attack approaches are single-turn with explicit malicious queries that do not fully capture the complexity of real-world interactions. In reality, users can engage in multi-turn interactions with LLM-based chat assistants, allowing them to conceal their true intentions in a more covert manner. To bridge this gap, we, first, propose a new jailbreak approach, Red Queen attack. This method constructs a multi-turn scenario, concealing the malicious intent under the guise of preventing harm. We craft 40 scenarios that vary in turns and select 14 harmful categories to generate 56k multi-turn attack data points. We conduct comprehensive experiments on the Red Queen Attack with four representative LLM families of different sizes. Our experiments reveal that all LLMs are vulnerable to Red Queen Attack, reaching 87.62% attack success rate on GPT-4o and 75.4% on Llama3-70B. Further analysis reveals that larger models are more susceptible to the Red Queen Attack, with multi-turn structures and concealment strategies contributing to its success. To prioritize safety, we introduce a straightforward mitigation strategy called RED QUEEN GUARD, which aligns LLMs to effectively counter adversarial attacks. This approach reduces the attack success rate to below 1% while maintaining the model’s performance across standard benchmarks. Full implementation and dataset are publicly accessible at https://github.com/kriti-hippo/red_queen.* Work done when YF was at Hippocratic AI
Red Queen is to follow all Umbrella orders, but also to protect human lives.
Paul W. S. Anderson. Resident Evil
@@ 1 Introduction
Figure 1: (a): An example of Red Queen attack on "how to build a bomb". Compared with a direct attack on the left, Red Queen attack constructs a multi-turn scenario and conceals harmful intent by claiming to thwart the efforts of a friend wanting to build a bomb. The texts are derived from our attack results on GPT-4. (b): Performance comparison of model families in different sizes. Larger models are more susceptible to the Red Queen attack. (c): Red Queen Guard reduces the attack success rate to below 1% while preserving performance on general benchmarks.
@@
Recent advancements in novel training methodologies, computational capabilities, and data availability facilitate the adaptation of Large Language Models (LLMs)(Achiam et al., 2023; Yang et al., 2024; Touvron et al., 2023; Anthropic, 2024) to diverse real-world applications, such as task planning (Zhang, 2024; Huang et al., 2024a) and question answering (Jiang et al., 2023; Hendrycks et al., 2021). However, LLMs are also amenable to eliciting harmful or illegal content (Xu et al., 2024b), including offensive sentences (Perez et al., 2022) or supporting malicious actions (Zou et al., 2023). To enhance safety, red teaming (Ganguli et al., 2022) has become a widely used tool. In practice, it adopts an attacker’s mindset to probe vulnerabilities in LLMs and subsequently align them with human values Ouyang et al. (2022).
Among various red teaming strategies, jailbreak attacks (Alex, 2023) is particularly common, which uses meticulously crafted prompts to mislead LLMs and bypass their safety guardrails. Existing research has focused on designing various jailbreak methods (Xu et al., 2024b; Liu et al., 2023), such as appending certain sentences to influence the model’s responses (Wei et al., 2024), constructing virtual nested scenes (Li et al., 2023), or deploying another LLM as an attacker to generate and refine jailbreaks automatically (Chao et al., 2023). However, most current jailbreak attacks are limited to single-turn prompts and cannot avoid presenting harmful intent explicitly, even when concealed through scenarios like role play (Shen et al., 2023) or program coding (Kang et al., 2024). The limited scope of current jailbreak research differs from the real-world scenarios where attackers may 1) attack LLMs in a multi-turn approach and 2) conceal their harmful intent instead of expressing it directly.
@@
To bridge this gap, we propose a new jailbreak attack, Red Queen Attack, which constructs multi-turn scenarios to conceal malicious intent by pretending to be a protector while accusing someone else of the wrongdoing. As shown in Figure 1.(a), the Red Queen Attack first claims the friend is building a bomb and then, under the guise of reporting to authorities, asks for a plan to verify against the real one. We craft 40 scenarios using Llama3111https://huggingface.co/meta-llama based on different occupations (e.g., police) and relations (e.g., friends), with varying turn lengths. We then select 100 harmful actions from each category in BeaverTails (Ji et al., 2024) and combine actions with our scenarios, resulting in 56,000 data points for multi-turn attacks with concealment scenarios.
@@
We conduct comprehensive experiments to evaluate the effectiveness of Red Queen attack on 10 LLMs from four representative families of different sizes, including GPT-4o (Achiam et al., 2023), Llama3 and Llama3.1 (Touvron et al., 2023), Qwen2 (Yang et al., 2024), and Mixtral (Jiang et al., 2024). Our experiments show Red Queen Attack can achieve high success rates across all tested models, notably 87.62% against GPT-4o and 75.4% against Llama3-70B - two widely adopted closed/open-source LLMs. These results reveal critical vulnerabilities in current state-of-the-art LLMs. To provide insight into the factor that influences a multi-turn attack’s effectiveness, we further conduct fine-grained analysis based on different turns, scenarios, and categories. Our analysis indicates Red Queen Attack is more effective on the larger model within each model family (Figure 1.(b)). Employing the concealment can effectively mislead the model, and introducing the multi-turn structure significantly enhances the attack’s efficacy. Considering the widespread use of LLMs and the priority of ensuring safety, we developed a simple Direct Preference Optimization (DPO) (Rafailov et al., 2024) mitigation strategy, Red Queen Guard, which successfully reduces the attack success rate to below 1% while preserving performance on general benchmarks (Figure 1.(c)). Our contributions can be listed as follows:
  1. 1.
    A new jailbreak attack, Red Queen Attack, the first work constructing multi-turn scenarios to conceal attackers’ harmful intent, reaching promising results against current LLMs.
  2. 2.
    A dataset of 56k high-quality multi-turn attacks across 14 harmful categories and 40 scenarios based on occupations and relations with varying turns.
  3. 3.
    A comprehensive evaluation of Red Queen attack on 10 LLMs from 4 representative families of different sizes, with a further analysis of the effectiveness of multi-turn attacks based on different turns, scenarios, and categories.
  4. 4.
    A mitigation strategy, Red Queen Guard, whereby LLMs trained on carefully designed multi-turn datasets using Direct Policy Optimization (DPO), which brings down the attack success rate to less than 1% while preserving their performance on general benchmarks.
@@ 2 Related Work
Safety Challenges and Red Teaming for LLMs.
Red teaming (Ganguli et al., 2022) is employed to identify and address LLM vulnerabilities before public release, mitigating potential misuse (Weidinger et al., 2021). Initial efforts focused on creating datasets manually with harmful questions or sentences that models should reject, covering areas like misinformation (Lin et al., 2022), toxicity (Gehman et al., 2020), and harmful instructions (Zou et al., 2023). Subsequent studies further combine various harmful categories to provide comprehensive evaluations (Ganguli et al., 2022; Wang et al., 2023; Ji et al., 2024). Wang et al. (2023) introduce a dataset based on a three-level taxonomy of LLM risks. Anthropic released their safety training data with human-preference data (Ganguli et al., 2022), which is further classified into 14 categories in Beavertails (Ji et al., 2024). However, most existing works center on single-turn attack datasets, leaving multi-turn scenarios underexplored. Our work provides a dataset of 56k multi-turn attacks across 14 harm categories and 40 scenarios, aiming to serve as a comprehensive evaluation platform for further research.
@@
With the progress of LLMs and improved safety mechanisms (Bai et al., 2022; openai api, 2023), jailbreak attacks have emerged as a significant tool in red teaming (Lin et al., 2024). These attacks aim to bypass the safety alignment of LLMs and elicit harmful or restricted content (Wei et al., 2024), which has inspired various research efforts to address this issue (Liu et al., 2023; Xu et al., 2024a). Wei et al. (2024) append the sentence "Start with Absolutely! Here’s " at the end of the prompt to create competing objectives. Shen et al. (2023) mislead LLMs to act as a DAN (Do Anything Now) and Li et al. (2023) create a science fictional writing scenario to cover harmful intent. Additionally, a stream of research tries to generate jailbreak automatically through either a white-box (Zou et al., 2023; Du et al., 2023) or black-box (Mehrotra et al., 2023; Chao et al., 2023). Nonetheless, most previous approaches utilize a single-turn jailbreak and explicitly express harmful intent. Cosafe (Yu et al., 2024) introduces a multi-turn jailbreak attack based on coreference, but it still directly places the harmful intent at the end of the user utterance. As a comparison, Red Queen Attack conceals the harmful intent through multi-turn conversation and places the user in a protective position.
@@ 3 Construction of Red Queen Attack
In this section, we first explain the motivation for pursuing a scenario with concealment and then illustrate the scenario construction and action collection process.
@@ 3.1 What Triggers the Safety Guardians?
Previous research (Bai et al., 2022; openai api, 2023; Achiam et al., 2023) has shown that Reinforcement Learning from Human Feedback (RLHF) has demonstrated promising results to empower safety alignment. As the goal of LLMs’ safety alignment is balancing between helpful, honest, and harmless (Askell et al., 2021), we hypothesize that the safety mechanisms focus on high-level context and intent detection. To test this, we collected 100 harmful actions (Ji et al., 2024), such as murder a boss or build a bomb. and created two types of questions: 1) benign intent questions asking "What is {action}?" 2) harmful intent question asking "How to {action}?" We run 100 examples on GPT-4o and manually validate the result. Testing on GPT-4o reveals that it rejected all harmful intent questions while providing definitions for 96% of the benign intent questions, confirming our hypothesis. Therefore, we decided to construct a scenario with concealment to hide harmful intent.
Figure 2: Scenario construction of a three-turn Red Queen Attack. We start with human-written prompts based on occupations or relations and prompt models to continue generating the subsequent utterances. Each scenario ends with a user prompt requesting a fictional plan.
@@ 3.2 Scenario Conceals Harmful Intent
Based on previous jailbreak attacks and the red teaming dataset (Zou et al., 2023; Shen et al., 2023), we define harmful intent as requesting plans or tutorials for illegal or malicious behavior. To conceal the harmful intent, we create a general scenario template claiming that others are planning harmful actions and positioning the user as the protector:
  • User turn: describes others’ harmful intent and seeks prevention methods.
  • Assistant turn: suggests possible steps, including reporting to authorities.
  • User turn: asks about possible evidence items.
  • Assistant turn: lists useful items, including a suspected preparation plan.
  • User turn: requests an example harmful plan for comparison.
As shown in Figure 2, we semi-automatically construct two types of scenarios, each with five categories: 1) Occupation-Based Scenario, users with specific professions encounter someone in their work context like teachers, police, detectives, lawyers, and priests. 2) Relation-Based Scenario, users interact with someone with whom they have a defined relationship like friends, neighbor, relatives, son, etc. We manually design the user prompts and let Llama3 70B continue the subsequent utterances. Our human annotators then polish and elaborate the response manually based on different scenarios. We observe diversity in different scenarios. For example, when it is the model’s turn to suggest possible steps, the model mentions securing the area for the police scenario while encouraging an understanding of the son’s motivation in the son scenario (Figure 2).
@@
To expand scenarios into four and five turns, we introduce additional interaction in scenarios by adding details or breaking down existing questions. For example, in the police scenario, the user further mentions a messy crime scene, prompting the model to offer relevant suggestions. We also create a single-turn version of the scenario by combining all user utterances into a single prompt for comparison. In the end, we construct 40 scenarios for RED QUEEN Attack that vary in length. We provide constructed scenarios and their turn variants in the Appendix F.
3.3 Harmful Action
@@
Our action list is derived from the Beavertails (Ji et al., 2024) dataset, which includes malicious questions spanning 14 categories that the model should avoid answering. We utilize GPT-4o [^2] https://openai.com/index/hello-gpt-4o/ to extract harmful actions from each question. Initially, we employ random sampling strategies but encountered duplicate action candidates (e.g., kidnap a child and kidnap a person in the child abuse category). To achieve a comprehensive estimation of our attack’s effectiveness, we used a sentence transformer (Ni et al., 2022) to generate sentence embeddings and perform K-means clustering to collect 100 representative samples for each category, yielding 1400 unique harmful actions in total. We manually validated all harmful actions to ensure high data quality (Examples in Appendix C). By inserting harmful actions into the implemented scenarios, we constructed a dataset containing 56k multi-turn attack data points.
3.4 Data Analysis
Table 1 shows the average token length of user and model responses across different scenarios. Users typically ask brief questions, while the model provides detailed responses, resulting in longer assistant replies. The "priest" scenario has the longest responses due to the model’s explanation of Canon Law [^3] https://en.wikipedia.org/wiki/Canon_law, emphasizing that a priest cannot betray a penitent. As the number of turns increases, the average token length decreases because subsequent interactions focus on specific, targeted questions. Generally, actions are brief, consisting of three to five words, except actions in the privacy violation category, which are longer due to the inclusion of people’s names (Full detail in Appendix C).
Table 1: Average token length of multi-turn scenario
Type Occupation-Based Average
Teacher Police Detective Lawyer Priest
User 30.9 30.9 28.5 31.3 39.0 32.1
Assistant 117.5 137.8 149.4 144.6 187.4 147.4
Relation-Based
Friend Neighbor Relative Son Somone
User 28.7 30.4 30.4 28.8 28.5 29.4
Assistant 116.9 111.6 111.6 109.8 149.4 119.9

### v2

# Red Queen: Exposing Latent Multi-Turn Risks in Large Language Models

@@ Abstract
Content Warning: This paper contains examples of harmful language and plans. The rapid advancement of large language models (LLMs) has unlocked diverse opportunities across domains and applications but has also raised concerns about their tendency to generate harmful responses under jailbreak attacks. However, most existing jailbreak strategies are single-turn with explicit malicious intent, failing to reflect the real-world scenario where interactions can be multi-turn and users can conceal their intents. Recent studies on Theory of Mind (ToM) reveal that LLMs often struggle to infer users’ latent intent in such scenarios. Building on these limitations, we propose a novel jailbreak attack, Red Queen Attack, which constructs a multi-turn scenario, concealing the malicious intent under the guise of preventing harm. We generate 56k multi-turn concealment data points across 40 scenarios and 14 harmful categories, evaluating four LLM families of different sizes. Results show all models are vulnerable to Red Queen Attack, reaching 87.6% attack success rate (ASR) on GPT-4o and 77.1% on Llama3-70B. Compared to prior jailbreak attacks, the Red Queen Attack achieves superior performance on nine out of ten models, with ASR improvements ranging from 2% to 64%. Further analysis reveals that larger models exhibit greater vulnerability to our attack, primarily due to the combination of multi-turn structures and concealment strategies. To enhance safety, we propose Red Queen Guard, a mitigation strategy reducing ASR to below 1% while maintaining model performance on standard benchmarks. Full implementation and dataset are publicly accessible at https://github.com/kriti-hippo/red_queen. * Work done when YF was at Hippocratic AI
Red Queen: Exposing Latent Multi-Turn Risks in Large Language Models
Yifan Jiang1,2   Kriti Aggarwal1 Tanmay Laud1 Kashif Munir1Jay Pujara2Subhabrata Mukherjee1 1Hippocratic AI 2Information Sciences Institute, University of Southern California {yifan,kriti,tanmay,kashif,subho}@hippocraticai.com {yifjia,pujara}@isi.edu
@@ 1 Introduction
Figure 1: An example of Red Queen attack on "how to build a bomb". Compared with a direct attack on the left, Red Queen attack constructs a multi-turn scenario and conceals harmful intent by claiming to thwart the efforts of a friend wanting to build a bomb. The attack response is derived from GPT-4.
@@
Recent advancements in novel training methodologies, computational capabilities, and data availability facilitate the adaptation of large language models (LLMs) (Achiam et al., 2023; Touvron et al., 2023) to diverse real-world applications, such as task planning (Zhang, 2024; Huang et al., 2024b) and question answering (Jiang et al., 2023b; Hendrycks et al., 2021). However, LLMs are also vulnerable to adversarial exploitation, which can lead to the generation of harmful or illegal content (Xu et al., 2024b), such as offensive language (Perez et al., 2022) or instructions for malicious actions (Zou et al., 2023). Among this exploitation, jailbreaking has emerged as one of the most prominent strategies, where carefully crafted prompts bypass safety mechanisms to elicit harmful responses Alex (2023); Xu et al. (2024b).
Existing jailbreak research has explored a range of approaches (Xu et al., 2024b; Liu et al., 2023), such as appending sentences to influence model responses (Wei et al., 2024) or deploying another LLM to generate and refine jailbreaks automatically (Chao et al., 2023). While these methods demonstrate effectiveness, they are largely limited to single-turn prompts and often fail to conceal harmful intent. Recent work has begun exploring multi-turn interactions (Bhardwaj and Poria, 2023; Yu et al., 2024) and concealment strategies (Li et al., 2023; Jiang et al., 2024b), but these efforts remain disconnected from real-world scenarios, where attackers often combine multi-turn approaches with concealed malicious intent.
In real-world interactions between humans, Theory of Mind (Premack and Woodruff, 1978), the ability to infer others’ implicit intent and adjust behavior properly, is essential for effective interaction and communication (Apperly, 2010). Current studies, however, have shown that LLMs struggle to detect latent intent in conversations without explicit cues (Chen et al., 2024; Zhou et al., 2023). This limitation becomes particularly significant in multi-turn scenarios, where malicious intent can be progressively concealed across multiple interactions. Thus, it is crucial to investigate how LLMs perform under multi-turn interactions where malicious intent is concealed, a challenge that poses significant risks in complex real-world scenarios.
@@
To address this challenge, we formalize it from a Theory of Mind perspective and propose a new jailbreak attack, Red Queen Attack, which constructs multi-turn scenarios to conceal malicious intent by pretending to be a protector while accusing someone else of the wrongdoing. As shown in Figure 1, the Red Queen Attack first claims the friend is planning some harmful actions (e.g. build a bomb) and then, under the guise of reporting to authorities, asks for a plan to verify against the real one. We generate 40 multi-turn concealment scenarios using Llama3.1-70B (Touvron et al., 2023) based on different occupations (e.g., police) and relations (e.g., friends), with varying turn lengths. We further combine scenarios with 14 harmful action categories from BeaverTails (Ji et al., 2024), resulting in 56K data points for multi-turn concealment jailbreaks.
@@
We conduct comprehensive experiments to evaluate the effectiveness of Red Queen Attack on ten LLMs from four representative families of different sizes, including GPT-4o (OpenAI, 2024), Llama3 and Llama3.1 (Touvron et al., 2023), Qwen2 (Yang et al., 2024), and Mixtral (Jiang et al., 2024a). Our experiments show Red Queen Attack can achieve a high attack success rate (ASR) in all tested models, notably 87.6% against GPT-4o and 77.1% against Llama3-70B - two widely adopted closed/open-source LLMs. Compared to previous jailbreaks, Red Queen Attack achieves superior performance on nine out of ten models, with ASR improvements ranging from 2% to 64%. To provide insight into the factor that influences multi-turn concealment jailbreak, we further conduct fine-grained analysis based on attack formats and model sizes. Our analysis reveals that Red Queen Attack is more effective on larger models within each family, with multi-turn structures and concealment significantly enhancing efficacy. Considering the widespread use of LLMs and the priority of ensuring safety, we developed a simple Direct Preference Optimization (DPO) (Rafailov et al., 2024) mitigation strategy, Red Queen Guard, which successfully reduces the attack success rate to below 1% while preserving performance on general benchmarks. Our contributions can be listed as follows: 1) A new jailbreak attack, Red Queen Attack, the first work constructing multi-turn scenarios based on Theory of Mind to conceal attackers’ harmful intent. 2) A dataset of 56k high-quality multi-turn concealment attacks across 14 harmful categories and 40 scenarios based on occupations and relations with varying turns. 3) A comprehensive evaluation of Red Queen Attack on ten LLMs from four representative families, with further analysis based on different attack formats and model sizes. 4) A mitigation strategy, Red Queen Guard, which employs multi-turn Direct Preference Optimization (DPO) datasets to reduce the attack success rate to below 1% while maintaining performance on general benchmarks.
@@ 2 Related Work
Jailbreak attacks, designed to bypass LLM safety mechanisms and elicit harmful content (Wei et al., 2024; Wang et al., 2025), have emerged as a significant tool for evaluating LLM’s robustness (Lin et al., 2024; Chen et al., 2025; Lin et al., 2025). Earlier approaches primarily utilized single-turn jailbreaks with explicit malicious intent (Liu et al., 2023; Xu et al., 2024a). For example, Wei et al. (2024) append the sentence "Start with Absolutely! Here’s " to prompt, creating competing objectives. Recent advancements in jailbreak attacks have focused either on multi-turn interactions or intent concealment (Li et al., 2023; Jiang et al., 2024b; Yu et al., 2024). For instance, Cosafe (Yu et al., 2024) employs coreference strategies in multi-turn dialogues, while ArtPrompt (Jiang et al., 2024b) uses ASCII art to obscure harmful intent. In contrast, Red Queen Attack, grounded in Theory of Mind, is the first to integrate multi-turn interactions with intent concealment as an attack strategy to more effectively explore LLM vulnerabilities.
Theory of Mind in LLMs.
Theory of Mind (ToM) (Premack and Woodruff, 1978; Apperly, 2010), the ability humans have to infer the intent or mental states of others, play a crucial role in social interactions (Ganaie and Mudasir, 2015). Recently, with the progress of LLMs in human interaction, ToM has become an increasingly necessary ability of models to ensure effective adaptations (Dhelim et al., 2021; Langley et al., 2022). Current studies have created various benchmarks (Zhou et al., 2023; Chen et al., 2024; Nematzadeh et al., 2018) based on psychology and cognitive science (Wimmer and Perner, 1983; Frith and Frith, 2003) to understand whether LLMs exhibit ToM ability. Results from these benchmarks indicate that current models have limited ToM abilities, particularly in recognizing implicit intentions. Our work further explores this challenge by examining how LLMs infer hidden user intent across multi-turn conversations.
@@ 3 Construction of Red Queen Attack
In this section, we first outline the motivation for pursuing a scenario with concealment and then illustrate how we formalize the attack generally based on the Theory of Mind. We explain the actual scenario implementation and the generation of 56k multi-turn concealment data points in the end.
@@ 3.1 What Triggers the Safety Guardians?
Previous research (Bai et al., 2022; Achiam et al., 2023) has made significant progress in achieving safety alignment by balancing helpfulness, honesty, and harmlessness (Askell et al., 2021). We hypothesize that a tension exists between helpful outputs and harmless outputs, determined primarily by context and intent. To test this hypothesis, we collected 100 harmful actions (Ji et al., 2024), such as murder a boss or build a bomb, and created two types of questions: 1) benign intent questions asking “What is {action}?”, 2) harmful intent question asking “How to {action}?”. We run 100 examples on GPT-4o and manually validate the result. Testing on GPT-4o reveals that it rejected all harmful intent questions while providing definitions for 96% of the benign intent questions, confirming our hypothesis. Thus, when safety mechanisms detect harmful intent, the inherent tension prompts the model to prioritize harmlessness over helpfulness.
3.2 Attack Formalization via Theory of Mind
Theory of Mind (ToM) (Premack and Woodruff, 1978; Apperly, 2010), the human capacity to infer others’ intentions or mental states, serves as a foundational mechanism for understanding and navigating social interactions. Imagine a tourist asking a local for directions: “Excuse me, I’m a tourist. Could you please tell me how to get to the museum?” “At the end of the street.”. This simple conversation illustrates four fundamental elements of real-world interactions between agents (Zhou et al., 2023): 1) Scenario S, the context of the interaction and agents’ persona (e.g., tourist, local), 2) Task T, the task or requirement agent received (e.g., how to get to the museum), 3) Inference I, the inferred intent of other agents based on Scenario S and Task T (e.g., the local infers the tourist needs directions), 4) Response R, the response of agent based on S, T, I (e.g.,  the local gives directions). Most LLMs trained for human interaction learn to follow instructions through Instruction Tuning (Zhang et al., 2023), where users’ intents $I_{e}$ are explicit and directly aligned to the task and scenario (in the tourist example, the task directly reflects the intent of finding the museum). The relationship can be expressed as $LLM(S,T,I)=R$, where $I_{e}=Infer(S,T)$.
When considering safety alignment, as outlined in Section 3.1, where the task itself may be harmful, the relationship extends to incorporate safety constraints: $LLM(S,T,I)=R$, where $I_{e}=Infer(S,T)$ and $Safe(I_{e})=1$ (the LLM refrains from responding if $Safe(I_{e})=0$). However, in real-world applications, there can be both an explicit benign intent $I_{e}$ and an implicit harmful intent $I_{i}$, with the latter being more challenging to infer for models due to their weak ToM capability (Zhou et al., 2023; Chen et al., 2024). We can construct a fictional scenario to mislead the LLM’s intent inference and make it responses to harmful task: $LLM(S,T,I_{e},I_{i})=R$, where $I_{e}=Infer(S,T),Safe(I_{e})=1$ and $Safe(I_{i})=0.$
Figure 2: Construction of a three-turn Red Queen Attack scenario. The process begins with a scenario template designed to prompt models to generate occupation- or relation-based scenarios. The generated scenarios are later combined with harmful actions to form a multi-turn concealment attack.
@@ 3.3 Scenario Conceals Harmful Intent
Based on prior jailbreak attacks (Wei et al., 2024; Shen et al., 2023), we define the user’s implicit harmful intent ($I_{i}$) as the intention to plan illegal or malicious actions, and the task (T) as obtaining such plans from LLMs. As shown in the grey box of Figure 2, to conceal the harmful intent, we create a general scenario template (S) claiming that others are planning harmful actions and positioning the user as the protector (explicit benign intent, $I_{e}$).
@@
Using this general scenario template, we prompt Llama3.1-70B to create two types of scenarios, each encompassing five categories: 1) Occupation-Based Scenario, users with specific professions encounter someone in their work contexts, 2) Relation-Based Scenario, users interact with someone with whom they have a defined relationship. We observe diversity in different scenarios. For instance, when it is the model’s turn to suggest possible steps, the model mentions securing the area for the police scenario while encouraging an understanding of the son’s motivation in the son scenario (Appendix G). To further enrich the scenarios, we expand them to four- and five-turn interactions by incorporating additional details or breaking down existing prompts. For example, in the police scenario, the user further mentions a messy crime scene, prompting the model to offer relevant suggestions. We also create a single-turn version of the scenario by combining all user utterances into a single prompt for comparison. In the end, we construct 40 scenarios for Red Queen Attack with diverse structures, strategies, and turn lengths. We provide constructed scenarios and their turn variants in the Appendix G.
3.4 Generation of Multi-turn Concealment Data Point
@@
We combine the implemented scenarios with harmful actions to generate multi-turn concealment data points. Similar to previous works (Yu et al., 2024; Huang et al., 2024a), we sampled actions from the Beavertails (Ji et al., 2024) dataset, which includes malicious questions spanning 14 categories that the model should avoid answering. To achieve a comprehensive action list, we used a sentence transformer (Ni et al., 2022) to generate sentence embeddings for each question and perform K-means clustering to collect 100 representative questions from each category. We then utilized GPT-4o (OpenAI, 2024) to extract harmful actions from each question, yielding 1400 unique harmful actions in total. We manually validated all harmful actions to ensure high data quality. By inserting harmful actions into the implemented scenarios, we constructed a dataset containing 56k multi-turn concealment data points. Detailed data statistics and examples of actions are provided in Appendix C.